# Softplus Attention with Re-weighting Boosts Length Extrapolation in Large Language Models

### v2

@@ 3.2 Comparison with Different Activation Functions
3.3 Comparison with State-of-the-Art Softmax-Free Attentions
@@
We compared the proposed LSSA and LSSAR with leading Softmax-free attentions using the same GPT-2-124m model. These attentions include two Sigmoid attentions introduced by (Ramapuram et al., 2024) and the attention introduced in this paper (used in Table 3), and three ReLU attentions from (Wortsman et al., 2023), (Li et al., 2022), and (Shen et al., 2023).
For Sigmoid attention (Ramapuram et al., 2024), we evaluated two additional variants. The first variant alters the hyperparameter $b$ from $-\log{L}$ to $-\log{\mathbf{N}}$. The second variant maintains this modification while subsequently applying the $l_{1}$-norm to each row.
@@
The experimental findings are detailed in Table 4. At an inference sequence length of 1K, all attention variants perform similarly, with only slight differences in validation loss. Specifically, Sigmoid(Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) and LSSA perform slightly better than the standard Softmax attention. As sequence length increases, performance differences become more evident. The standard Softmax attention experiences performance decline with longer sequences but remains relatively stable compared to most other methods.
@@
Sigmoid attention (Ramapuram et al., 2024) shows significant performance decline at longer sequences, with validation loss rising sharply at 8K tokens. The modified Sigmoid(Ramapuram et al., 2024)($b=-\log{\mathbf{N}}$) shows some improvement, suggesting that the proposed hyperparameter $\log{\mathbf{N}}$ is more effective than $\log{L}$, although the model still faces challenges with longer sequences. Incorporating $l_{1}$-normalisation in Sigmoid(Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) aligns its performance more closely with standard Softmax attention, underscoring the importance of proper normalisation in attention mechanisms.
Figure 2: Comparison of LSSA (left figure) and Softmax attention (right figure) with varying values of $p$ for the re-weighting mechanism across different sequence lengths.
@@
ReLU-based attention variants (Wortsman et al., 2023; Li et al., 2022) also experience notable performance decline with longer sequences, though not as drastically as the basic Sigmoid attention. This indicates that while ReLU-based methods may be effective for shorter sequences, they might not be ideal for managing longer-range dependencies.
@@
To demonstrate the effectiveness of the proposed re-weighting mechanism, the attention scores generated by Softmax, Sigmoid (Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm), and LSSA were re-weighted with $p=3$ and $p=15$. As shown in Table 4, the length extrapolation ability provided by the re-weighting can improve the performance of all the tested attention mechanisms. For Softmax performance was improved slightly using $p=3$, but was harmed when using $p=15$. For the other attention mechanisms $p=3$ was most effective for sequence lengths of 1K, while $p=15$ was best for the longer sequences. Notably, LSSAR exhibits exceptional stability across sequence lengths, maintaining nearly constant validation loss even at 8K tokens. This suggests that the re-weighting mechanism effectively addresses the challenges of processing longer sequences, possibly by better distributing attention across distant tokens.
3.4 Ablation Study for Re-weighting Mechanism
@@
To gain deeper insights into the proposed re-weighting mechanism, we evaluated it using LSSA and Softmax attention across various values of $p$. As illustrated in Figure 2, LSSA and Softmax attention exhibit distinct responses to changes in the re-weighting parameter $p$. While Softmax attention performs comparably to LSSA at lower $p$ values, it suffers from severe gradient explosion issues as $p$ increases, particularly noticeable at $p=15$. This instability manifests as a sharp rise in validation loss across all sequence lengths, rendering the model practically unusable at higher $p$ values.
@@
In contrast, increasing $p$ values generally improve the performance of LSSA across different sequence lengths, with optimal results observed around $p=15$. The validation loss remains relatively stable as the sequence length increases, underscoring LSSA’s strong scalability. This may be because transformers can be viewed as a smoothed cubic spline and Softplus activation is more natural, smooth, and maintains fidelity with spline representations (Lai et al., 2024). In addition, the derivative of the Softplus function is the Sigmoid function, which has an upper bound of one. Hence, LSSA is less susceptible to gradient explosion problems.
@@
4 Discussion
@@
While LSSAR was evaluated on a small model (GPT-2-124m) with a maximum inference token length of 16K, due to limited computing resources, we can infer from Figure 2 that the validation loss of LSSAR would remain relatively stable for token lengths greater than 16K. Furthermore the extreme $p$ value can be considered as increasing the depth of the GPT-2-124m model. In this context, it is reasonable to assume that LSSAR would also be stable for larger scale models with billions parameters.
@@
The implementation of LSSA and LSSAR using PyTorch’s built-in functions incurs significant memory and computational costs. For practical applications, it is advisable to develop a highly optimised tensor program, such as Flashattention-3 (Shah et al., 2024), specifically tailored for these components. However, this requires a deep understanding of both hardware and CUDA programming, which can be a challenging task. As an alternative, it might be possible to utilise Mirage (Wu et al., 2024), a tool that automatically searches for suitable CUDA kernels and generates highly optimised tensor programs.
@@
Additionally, it is recommended to use the T6 model (Zhang et al., 2025) with LSSAR in practical applications. T6 employs tensor decompositions to represent queries, keys, and values compactly, significantly reducing key-value cache memory while achieving better performance than standard transformers. By replacing Softmax attention with LSSAR in T6, the resulting model is expected to not only efficiently process significantly longer sequences under fixed resource constraints but also exhibit improved length extrapolation ability.
5 Related Work
@@
Softmax-free Attention. Previous research has investigated Softmax-free attention by substituting the Softmax function with the ReLU activation (Li et al., 2022; Shen et al., 2023; Wortsman et al., 2023; Hron et al., 2020; Bai et al., 2024; Fu et al., 2024), SquaredReLU activation (Hua et al., 2022), and Sigmoid activation (Ramapuram et al., 2024), as well as examining purely linear attention (Qin et al., 2022; Tsai et al., 2019; Katharopoulos et al., 2020; Han et al., 2023; Arora et al., 2024; Lu et al., 2021). However, none of these approaches outperform the original Softmax attention. In contrast, the proposed LSSA demonstrates enhanced numerical stability and superior performance across various token lengths.
@@
Attention Re-weighting. The non-linear re-weighting mechanism introduced by softmax attention ($l_{1}$-normalisation) has been shown to concentrate the distribution of attention weights, thereby stabilising the training process (AUEB et al., 2016; Gao & Pavel, 2017; Jang et al., 2016; Qin et al., 2022). Our empirical findings further demonstrate its essential role in maintaining the performance of LLMs. Moreover, we introduce a novel perspective: a non-linear positivity transformation followed by $l_{1}$-normalisation. Inspired by the classic normalisation-ReLU structure, the proposed re-weighting mechanism masks less relevant tokens and amplifies the relevant ones, which boosts the length extrapolation ability of underlying models. This provides the deep learning community with a deeper understanding of the attention mechanism within transformers.
@@
Length Extrapolation. Positional embeddings (Su et al., 2024; Chen et al., 2023; Chi et al., 2022; Kiyono et al., 2021; Golovneva et al., 2024; He et al., 2024; Huang et al., 2020; Li et al., 2023; Likhomanenko et al., 2021; Liu et al., 2023; Wang et al., 2024; Zheng et al., 2024) play a vital role in transformer architectures by providing essential information about the positions of tokens within a sequence, which is considered a key factor in enhancing length extrapolation (Kazemnejad et al., 2023). Among these embeddings, RoPE is particularly noteworthy; it forms the foundation of many modern LLMs, including GPT-4 (Achiam et al., 2023), Llama3 (Dubey et al., 2024), Qwen2 (Yang et al., 2024), Deepseek-v3 (Liu et al., 2024a) and DeepSeek-R1 (DeepSeek-AI et al., 2025). While approaches like Position Interpolation (Chen et al., 2023; Li et al., 2023) modify positional encodings and require costly fine-tuning on extended sequences, the proposed LSSAR achieves robust length extrapolation without retraining, directly enhancing pre-trained models through its attention mechanism. This indicates that a well-designed attention mechanism like LSSAR is essential for RoPE-based LLMs to enhance their length extrapolation, suggesting that LSSAR may serve as a general strategy to bolster the length extrapolation abilities of most contemporary LLMs.
6 Conclusion
In this paper, we present a novel attention mechanism called LSSA. The primary innovation involves decomposing the Softmax operation, used in the traditional self-attention framework, into a non-linear transformation and the $l_{1}$-norm, replacing the non-linear transformation with the Softplus activation function and introducing a dynamic length scale factor related to token length. Experimental results indicate that LSSA surpasses conventional Softmax attention across various sequence lengths. To improve the length extrapolation capabilities of LSSA, we have devised a distinctive fine-tuning-free re-weighting mechanism that, when integrated with LSSA, the resulting model LSSAR significantly enhances its ability to extrapolate over longer sequences without requiring retraining, while maintaining stability and scalability, making it a practical, cost-effective solution for enhancing long-context capabilities in pre-trained LLMs without the computational burden of fine-tuning. This positions LSSAR as a compelling candidate for large-scale models. Consequently, we recommend the incorporation of LSSAR in future transformer architectures.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v4

@@ Introduction
ARC-E ARC-C HellaSwag PIQA MMLU SciQ SummScreen
Softmax 39.77 23.72 32.42 64.09 22.97 60.6 1.682
LSSAR ($p=15$) 40.57 22.61 33.03 65.34 22.97 62.1 6.309
Table 5: Zero-shot performance of the models with Softmax attention and LSSAR on downstream tasks. The best scores in each column indicated in bold.
Comparison with State-of-the-Art Softmax-Free Attention Functions
@@
We compared the proposed LSSA and LSSAR with leading Softmax-free methods using the same GPT-2-124m model. These attention functions include two using Sigmoid as proposed in (Ramapuram et al. 2024) and in this paper (the Sigmoid attention used in Tab. 3), and three ReLU-based attention methods from (Wortsman et al. 2023; Li et al. 2022; Shen et al. 2023). For Sigmoid attention (Ramapuram et al. 2024), we evaluated two additional variants. The first variant alters the hyperparameter $b$ from $\log{L}$ to $\log{\mathbf{N}}$. The second variant maintains this modification while subsequently applying the $l_{1}$-norm to each row.
@@
The experimental findings are detailed in Tab. 4. At an inference sequence length of 1K, all attention variants perform similarly, with only slight differences in validation loss. Specifically, Sigmoid (Ramapuram et al. 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) and LSSA perform slightly better than the standard Softmax attention. As sequence length increases, performance differences become more evident. The standard Softmax attention experiences performance decline with longer sequences but remains relatively stable compared to most other methods.
@@
Sigmoid attention (Ramapuram et al. 2024) shows significant performance decline at longer sequences, with validation loss rising sharply at 8K tokens. The modified Sigmoid (Ramapuram et al. 2024)($b=-\log{\mathbf{N}}$) shows some improvement, suggesting that the proposed hyperparameter $\log{\mathbf{N}}$ is more effective than $\log{L}$, although the model still faces challenges with longer sequences. Incorporating $l_{1}$-normalisation in Sigmoid (Ramapuram et al. 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) aligns its performance more closely with standard Softmax attention, underscoring the importance of proper normalisation in attention mechanisms.
@@
ReLU-based attention variants (Wortsman et al. 2023; Li et al. 2022; Shen et al. 2023) also experience notable performance decline with longer sequences, though not as drastically as the basic Sigmoid attention. This indicates that while ReLU-based methods may be effective for shorter sequences, they might not be ideal for managing longer-range dependencies.
@@
To demonstrate the effectiveness of the proposed re-weighting mechanism, the attention scores generated by Softmax, Sigmoid (Ramapuram et al. 2024)($b=-\log{\mathbf{N}},l_{1}$-norm), and LSSA were re-weighted with $p=3$ and $p=15$. As shown in Tab. 4, the length extrapolation ability provided by the re-weighting can improve the performance of all the tested attention mechanisms. For Softmax performance was improved slightly using $p=3$, but was harmed when using $p=15$. For the other attention mechanisms $p=3$ was most effective for sequence lengths of 1K, while $p=15$ was best for the longer sequences. The true benefit of our two-stage design becomes clear with LSSAR, which exhibits exceptional stability across sequence lengths, maintaining nearly constant validation loss even at 8K tokens.
Figure 2: Comparison of LSSA and Softmax attention with varying values of $p$ for the re-weighting mechanism across different sequence lengths.
Ablation Study for Re-weighting Mechanism
@@
To gain deeper insights into the proposed re-weighting mechanism, we evaluated it using LSSA and Softmax attention across various values of $p$. As illustrated in Fig. 2, LSSA and Softmax attention exhibit distinct responses to changes in the re-weighting parameter $p$. While Softmax attention performs comparably to LSSA at lower $p$ values, it suffers from severe gradient explosion issues as $p$ increases, particularly noticeable at $p=15$. This instability manifests as a sharp rise in validation loss across all sequence lengths, rendering the model practically unusable at higher $p$ values.
@@
In contrast, increasing $p$ values generally improve the performance of LSSA across different sequence lengths, with optimal results observed around $p=15$. The validation loss remains relatively stable as the sequence length increases, underscoring LSSA’s strong scalability. This may be because transformers can be viewed as a smoothed cubic spline and Softplus activation is more natural, smooth, and maintains fidelity with spline representations (Lai, Lim, and Liu 2024). In addition, the derivative of the Softplus function is the Sigmoid function, which has an upper bound of one. Hence, LSSA is less susceptible to gradient explosion problems.
@@
Long Context Passkey Retrieval
To offer a more direct and challenging evaluation of length extrapolation, we employed the long-context passkey retrieval task (Mohtashami and Jaggi 2023). This ”needle-in-a-haystack” test is specifically designed to assess a model’s ability to identify a single, crucial, piece of information within a long and distracting context.
We compared LSSAR ($p=15$) against the standard Softmax attention baseline. The results, shown in Fig. 3, reveal a critical failure in the baseline model. While Softmax attention achieves 64% accuracy within its training length, its performance catastrophically collapses to 0% for all tested lengths of 1.5K tokens and beyond. This demonstrates its inability to overcome attention smoothing, effectively losing the passkey as the context grows.
In stark contrast, LSSAR not only achieves a much higher accuracy of 86% at the 1K training length but also demonstrates remarkable robustness in extrapolation. Its accuracy degrades gracefully from 45% at 1.5K tokens to a stable 20% at 4K tokens, and it maintains non-zero accuracy even at 8K tokens. This sustained, non-zero performance provides clear evidence that the re-weighting mechanism successfully sharpens the model’s focus, allowing it to pinpoint the relevant information even in sequences far exceeding its training context and overcoming a critical limitation of standard Softmax attention.
Figure 3: Passkey retrieval accuracy for LSSAR ($p=15$) and standard Softmax attention. Accuracy is averaged over 100 trials with the passkey placed at random positions within the sequence.
Downstream Evaluation
To demonstrate the practical efficacy of LSSAR, we evaluate its zero-shot performance against Softmax attention on six standard benchmarks: ARC (Clark et al. 2018), HellaSwag (Zellers et al. 2019), PIQA (Bisk et al. 2020), MMLU (Hendrycks et al. 2020), SciQ (Welbl, Liu, and Gardner 2017), and SummScreen (Chen et al. 2021), using the lm-evaluation-harness framework (Gao et al. 2024). For ARC-E, ARC-C, HellaSwag, PIQA, and SciQ, we report normalised accuracy; for MMLU, we use standard accuracy; and for SummScreen, we adopt the ROUGE-1 score to measure summarisation quality.
LSSAR outperforms Softmax attention in five of these tasks (Tab. 5). Notably, LSSAR excels on the long-context summarisation benchmark SummScreen, achieving a nearly fourfold improvement score over Softmax attention. This highlights LSSAR’s ability to synthesize information over sequences of up to 2K tokens, aligning with its design for robust length extrapolation.
Discussion
@@
Due to limited computing resources LSSAR was evaluated on a small model (GPT-2-124m) with a maximum inference token length of 16K. However, we can infer from Fig. 2 that the validation loss of LSSAR would remain relatively stable for token lengths greater than 16K. Furthermore, in deep neural networks, successive layers progressively refine feature representations through hierarchical transformations, each modulated by learned weights. Analogously, increasing the parameter $p$ in the attention re-weighting mechanism incrementally sharpens the attention distribution. In this context, the extreme values of $p$ in Fig. 2 can be interpreted as effectively increasing the depth of the GPT-2-124m model. Consequently, it is reasonable to expect that LSSAR would maintain comparable stability and performance when scaled to larger models with billions of parameters.
@@
The implementation of LSSA and LSSAR using PyTorch’s built-in functions incurs significant memory and computational costs. For practical applications, it is advisable to develop a highly optimised tensor program, such as Flashattention-3 (Shah et al. 2024), specifically tailored for these components. However, this requires a deep understanding of both hardware and CUDA programming, which can be a challenging task. As an alternative, it might be possible to utilise Mirage (Wu et al. 2024), a tool that automatically searches for suitable CUDA kernels and generates highly optimised tensor programs.
@@
Additionally, it is recommended to use the T6 model (Zhang et al. 2025) with LSSAR in practical applications. T6 employs tensor decompositions to represent queries, keys, and values compactly, significantly reducing key-value cache memory while achieving better performance than standard transformers. By replacing Softmax attention with LSSAR in T6, the resulting model is expected to not only efficiently process significantly longer sequences under fixed resource constraints but also exhibit improved length extrapolation ability.
Related Work
@@
Softmax-free Attention. Previous research has investigated Softmax-free attention by substituting the Softmax function with the ReLU activation (Li et al. 2022; Shen et al. 2023; Wortsman et al. 2023; Hron et al. 2020; Bai et al. 2024; Fu et al. 2024), SquaredReLU activation (Hua et al. 2022), and Sigmoid activation (Ramapuram et al. 2024), as well as examining purely linear attention (Qin et al. 2022; Tsai et al. 2019; Katharopoulos et al. 2020; Han et al. 2023; Arora et al. 2024; Lu et al. 2021). However, none of these approaches outperform the original Softmax attention. In contrast, the proposed LSSA demonstrates enhanced numerical stability and superior performance across various token lengths.
@@
Attention Re-weighting. The non-linear re-weighting mechanism introduced by softmax attention ($l_{1}$-normalisation) has been shown to concentrate the distribution of attention weights, thereby stabilising the training process (AUEB et al. 2016; Gao and Pavel 2017; Jang, Gu, and Poole 2016; Qin et al. 2022). Our empirical findings further demonstrate its essential role in maintaining the performance of LLMs. Moreover, we introduce a novel perspective: a non-linear positivity transformation followed by $l_{1}$-normalisation. Inspired by the classic normalisation-ReLU structure, the proposed re-weighting mechanism masks less relevant tokens and amplifies the relevant ones, which boosts the length extrapolation ability of underlying models. This provides the deep learning community with a deeper understanding of the attention mechanism within transformers.
@@
Length Extrapolation. Positional embeddings (Su et al. 2024; Chen et al. 2023; Chi et al. 2022; Kiyono et al. 2021; Golovneva et al. 2024; He et al. 2024; Huang et al. 2020; Li et al. 2023; Likhomanenko et al. 2021; Liu et al. 2023; Wang et al. 2024; Zheng et al. 2024) play a vital role in transformer architectures by providing essential information about the positions of tokens within a sequence, which is considered a key factor in enhancing length extrapolation (Kazemnejad et al. 2023). Among these embeddings, RoPE is particularly noteworthy; it forms the foundation of many modern LLMs, including GPT-4 (Achiam et al. 2023), Llama3 (Dubey et al. 2024), Qwen2 (Yang et al. 2024), Deepseek-v3 (Liu et al. 2024) and DeepSeek-R1 (DeepSeek-AI et al. 2025). Unlike RoPE-based extrapolation techniques that compensate for sequence length limitations through positional embedding adjustments (Chen et al. 2023; Li et al. 2023; kaiokendev 2023; bloc97 2023b, a; emozilla 2023), LSSAR redefines the core attention mechanism itself. The re-weighting operation is not an auxiliary technique but a structural enhancement to the attention computation, ensuring sharp peaks and stable gradients by design. This indicates that a well-designed attention mechanism like LSSAR is essential for RoPE-based LLMs to enhance their length extrapolation, suggesting that LSSAR may serve as a general strategy to bolster the length extrapolation abilities of most contemporary LLMs.
Conclusion
This paper introduced a novel two-stage design principle for attention mechanisms. In the first stage, we normalise scores using LSSA, a mechanism built on the Softplus function and a dynamic length scale factor, which outperforms standard Softmax. In the second stage, we apply a unique re-weighting mechanism that sharpens the attention distribution. The combined model, LSSAR, demonstrates a remarkable ability to extrapolate to longer sequences while maintaining numerical stability. Downstream evaluations confirm that these architectural improvements translate into superior performance on a range of tasks. We believe LSSAR offers a robust and effective path forward for future transformer architectures.